# How does the entanglement entropy of a many-body quantum system change after a single measurement?

### v3

@@ IV The quantum jump measurement protocol
Figure 4: The probability distribution of the change in entanglement entropy $\Delta S_{qj}$ Eq. (9), computed employing Eq. (7), due to a quantum jump in the QJ protocol, for sufficiently late times so that EE has reached the saturation value. The $P(\Delta S_{qj})$ is rescaled by multiplying the system size $L$. The tail is independent of system size after rescaling, signaling the tail is dominated by the boundary region. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 4 to 4.
In order to gain a more quantitative understanding of the distribution, especially this left tail, we investigate the spatial dependence of $\Delta S_{qj}$ Eq. (8), see Fig. 5. Interestingly, we find very different behavior for sites corresponding to the boundary between the two halves in which we split the system to compute the EE. For weak monitoring strength, the distribution of points at, and around, the subsystem boundary is broader, asymmetric, and with a much softer peak at zero. By contrast, for the rest of sites not very close to the boundary, the distribution has a peak at zero with a much narrower distribution. In the strong monitoring limit, and for sites far from the boundary, we observe a complete quantum Zeno effect $\Delta S_{qj}=0$ while in the region around the boundary, the distribution, though peaked at zero, has a well defined symmetric distribution. These observations indicate qualitative differences between the region around boundary and the rest of sites. The precise length for this boundary region may depend on the monitoring strength. Far from the boundary, the effect of measurements becomes weak in the sense that the change of EE is small even for relatively weak monitoring strengths.
Figure 5: $\Delta S_{qj}$ Eq. (9) as a function of the measured sites $r=1,\ldots,L$. The left plot is for the weak monitoring rate $\gamma=0.1$ and system size $L=1024$, while the right plot is for the strong monitoring rate $\gamma=3$ and the system size is $L=512$. Note the stark differences between boundary points, separating the two subsystems defining the EE, and the rest.
Therefore, we expect that when the subsystem size is much larger than this boundary region, $P(\Delta S_{qj})$ would be independent of the subsystem size $\ell$, which is illustrated in Fig. 6. In the weak monitoring limit $\gamma=0.1$, when the subsystem size is smaller than $128$, $P(\Delta S_{qj})$ changes substantially from a broader distribution to a narrower distribution and finally become insensitive to $\ell$. When the monitoring is stronger, $\gamma\geq 0.5$, the boundary region is much smaller than subsystem size $\ell\geq 16$ that we investigate, so $P(\Delta S_{qj})$ do not depend on the subsystem size. For the projective measurement protocol, we also obtain a similar results (not shown).
@@
An explicit calculation of the distribution function separating sites close to the boundary from those in the bulk, see Fig. 7, confirms the qualitative picture above. For weak monitoring, the distribution function for the boundary point is a Gaussian with a small peak at zero. As more sites around the boundary are added, the distribution develops a narrow exponential tail for $\Delta S_{qj}\gtrsim 0$. We note that rare instances of $\Delta S_{qj}\gtrsim 0$ occur due to the fact that the measurement not only disentangles the target site but also rearranges the remaining particles across the rest of the L1L-1 sites, effectively enhancing entanglement in unmeasured regions. This counterintuitive effect highlights the non-local feedback inherent in quantum measurement dynamics. For bulk sites, the distribution is asymmetric with exponential tails characterized by different decay rates. For strong monitoring, the distribution is effectively a delta function at zero for bulk points (note the scale of the axis), while for boundary points, though there is also a peak at zero, we observe an asymmetric distribution with a broad support that covers all possible values of $\Delta S_{qj}$.
@@
Figure 7: The distribution function of $\Delta S_{qj}$ Eq. (9) for specific sites. The upper panel: the monitoring rates is $\gamma=0.1$ and the system size is $L=1024$. The bottom panel: the monitoring rates is $\gamma=3.0$ and the system size is $L=512$. Here sites 1, 2 and 3 stand for sites around the boundary of the subsystem. Site 4 stands for the site at the center of the subsystem. Specifically, sites 1 are sites $[r=1,r=L/2,r=L/2+1,r=L]$, sites 2 are [r=1+5,r=L/25,r=L/2+6,r=L5][r=1+5,r=L/2-5,r=L/2+6,r=L-5], sites 3 are [r=1+15,r=L/215,r=L/2+16,r=L15][r=1+15,r=L/2-15,r=L/2+16,r=L-15] and sites 4 $[r=L/4,r=3L/4]$. Note that for Fig. 7, the scale of $\Delta S_{qj}$ is of the order $10^{-10}$ and the maximum is $10^{10}$ so it is effectively a Dirac delta function.
Figure 8: The density-map of $\Delta S_{qj}$ as a function of the occupation number of the selected site that are going to be measured. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 8 to 8. The system size for the upper panel is $L=1024$, and the lower panel is $L=512$.
@@
In the strong monitoring limit, $n_{i}$ can take any value though the change in EE has a cut-off value $\ln{2}\approx 0.69$, reflecting the maximal entanglement contribution from a single particle. As mentioned earlier, this maximum value can only occur at the subsystem boundary sites because, for the rest of sites, measurements induce little change in $\Delta S_{qj}$. Therefore, a measurement would only affect particles around the boundary.
@@
where $\tau$ is the time between quantum jumps and $\Delta S_{nH}=S_{A}(t+\tau)-S_{A}(t)$ computed from Eq. (7). Therefore, $\delta S_{nH}$ captures the system intrinsic capacity to regenerate entanglement between measurements. Since we are interested in the late time dynamics of entanglement for which the average EE has reached the saturation value and therefore does not have a net growth in time, we expect [68] that, on average, the generation of entanglement due to the non-Hermitian evolution will compensate its destruction due to measurements $\overline{\delta S_{nH}}\approx\overline{\Delta S_{qj}}/\bar{\tau}$, where $\overline{\penalty 10000\ \bullet\penalty 10000\ }$ stands for the average over quantum trajectories. Results depicted in the left plot of Fig. 9 confirm that this is the case for all considered monitoring strengths. We stress that this cancellation is a property of the average only. We shall see that the distribution of $\delta S_{nH}$ and other properties are qualitatively different from those due to quantum jumps $\Delta S_{qj}$.
@@
Figure 9: The average change in EE after a measurement as a function of the monitoring rate $\gamma$ using the 9 QJ protocol (left) and the 9 PM protocol (right). This is compared with the average speed of entanglement gain between measurements in the QJ protocol $\delta S_{nH}$ (left) and the PM protocol ($\delta S_{un}$) (right). As is expected, the creation of entanglement is compensated by the loss due to the measurement. The system size is $L=512$.
@@
Figure 10: The probability distribution of the change of EE $\delta S_{nH}$ Eq. (10), computed using Eq. (7), between measurements due to the non-hermitian evolution in the QJ protocol for different system sizes and times longer than the saturation time. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 10 to 10. The monitoring rate is the same as in Fig. 4.
@@
Figure 11: The entanglement change $\delta S_{nH}$ Eq. (10) due to the non-Hermitian evolution in the QJ protocol after measuring site $r$. The left plot is for the weak monitoring rate $\gamma=0.1$ and system size $L=1024$, while the right plot is for the strong monitoring rate $\gamma=3$. The system size is $L=512$. Different from the effect of quantum jumps $\Delta S_{qj}$, see Fig. 5, $\delta S_{nH}$ due to the non-Hermitian evolution does not show any dependence on the measuring site.
Figure 12: The density-map of $\delta S_{nH}$ for the QJ protocol as a function of the occupation number before the measurements. The monitoring rate are $\gamma=0.1$ (left), and $\gamma=3.0$ (right).
@@
Figure 13: The mutual information $I(d_{r})$ Eq. (11) between the subsystem AA and a site $M_{r}$ in the other subsystem as a function of the distance $d_{r}$ for the QJ protocol. The pink dashed line is a power-law fitting and the red dashed line is an exponential fitting. The system size is $L=512$.
@@ V The projective measurement protocol
Finally, we study the PM protocol corresponding to the standard picture in quantum mechanics where, after a measurement, the wavefunction collapses to one of the eigenstates of the Hermitian operator representing the observable being measured with a probability given by the Born’s rule. This protocol is similar to the quantum jump one. However, there are a few differences. Since there is no continuous monitoring, the evolution between measurements is Hermitian. Moreover, the outcome of the measurement is $n_{i}=0$ or $1$ while in the quantum jump protocol the only outcome is $n_{i}=1$. As a consequence, in the strong monitoring limit, we expect quantum Zeno effect around these two values.
Figure 14: The probability distribution $P(\Delta S_{pm})$ of the entanglement entropy change $\Delta S_{pm}$ Eq. (12), computed from Eq. (7) after saturation of the EE, due to a projective measurement of the occupation number (PM protocol). For convenience, the distribution $P(\Delta S_{pm})$ is multiplied by the system size $L$. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 14 to 14.
Figure 15: Density map of $\Delta S_{pm}$ Eq. (12) with the same parameters as in Fig. 8 (QJ protocol) but for the PM protocol. Results are similar for both protocols after taking into account that the potential outcomes in the PM protocol are $n_{i}=0,1$ while in the QJ protocol is only $n_{i}=1$.
Details of the protocol are given in Appendix A.3. As in the previous protocols, we also define the change in EE
@@
Figure 16: The density-map of $\delta S_{un}$ Eq. ( 12) for the PM protocol as a function of the occupation number right before the measurement. The monitoring rates are $\gamma=0.1$ (left), and $\gamma=3.0$ (right).
@@ VI Discussion and Conclusions
We have studied the distribution function of the change in EE after a measurement of the occupation number in a chain of free fermions using the QSD, QJ and PM protocols. QJ and PM protocols lead to asymmetric non-Gaussian distributions with a peak at zero that becomes stronger as the monitoring strength or the system size increases. More interestingly, for these two protocols, the distribution function is spatially inhomogeneous, namely, it is not the same for all sites. For boundary sites, separating the two sub-systems in the definition of the EE, the distribution in the weak monitoring limit is close to a Gaussian with a small peak at zero and a broad support. However, for the rest of sites, the support is one order of magnitude smaller and the distribution has asymmetric exponential tails. On average, the change in EE is negative, however, we have found that the distribution has an exponential tail for positive values so there is a finite probability that the EE increases as a result of a single measurement. It would be interesting a detailed characterization of the conditions for the existence of such measurements resulting in a positive change of the EE.
@@
The distribution for the QSD protocol is qualitatively different. For weak monitoring, it is Gaussian and size independent. As the monitoring strength increases, the distribution is still symmetric but deviations from Gaussianity are clearly observed which are well characterized by a distribution which is Gaussian around zero but with exponential tails. In the strong monitoring limit, the tail are still exponential but the Gaussian around zero turns into a strong peak suggesting a singularity at zero which is a signature of the Zeno effect.

### v4

@@ IV.1 Distribution of $\Delta S_{qj}$ after each measurement
Figure 4: The probability distribution of the change in entanglement entropy $\Delta S_{qj}$ in Eq. (9), computed employing Eq. (7), due to a quantum jump in the QJ protocol, for sufficiently late times so that EE has reached the saturation value. The $P(\Delta S_{qj})$ is rescaled by multiplying the system size $L$. The tail is independent of system size after rescaling, signaling the tail is dominated by the boundary region. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 4 to 4.
IV.2 Spatial dependence of $\Delta S_{qj}$
In order to gain a more quantitative understanding of the distribution, especially this left tail, we investigate the spatial dependence of $\Delta S_{qj}$ defined in Eq. (9), see Fig. 5. Interestingly, we find very different behavior for sites corresponding to the boundary between the two halves in which we split the system to compute the EE. For weak monitoring strength, the distribution of points at, and around, the subsystem boundary is broader, asymmetric, and with a much softer peak at zero. By contrast, for the rest of sites not very close to the boundary, the distribution has a peak at zero with a much narrower distribution. In the strong monitoring limit, and for sites far from the boundary, we observe a complete quantum Zeno effect $\Delta S_{qj}=0$ while in the region around the boundary, the distribution, though peaked at zero, has a well defined symmetric distribution. These observations indicate qualitative differences between the region around boundary and the rest of sites. The precise length for this boundary region may depend on the monitoring strength. Far from the boundary, the effect of measurements becomes weak in the sense that the change of EE is small even for relatively weak monitoring strengths.
Figure 5: $\Delta S_{qj}$ in Eq. (9) as a function of the measured sites $r=1,\ldots,L$. The color scale indicates the normalized density of measurement events. The left plot is for the weak monitoring rate $\gamma=0.1$ and system size $L=1024$, while the right plot is for the strong monitoring rate $\gamma=3$ and the system size is $L=512$. Note the stark differences between boundary points, separating the two subsystems defining the EE, and the rest.
Therefore, we expect that when the subsystem size is much larger than this boundary region, $P(\Delta S_{qj})$ would be independent of the subsystem size $\ell$, which is illustrated in Fig. 6. In the weak monitoring limit $\gamma=0.1$, when the subsystem size is smaller than $128$, $P(\Delta S_{qj})$ changes substantially from a broader distribution to a narrower distribution and finally become insensitive to $\ell$. When the monitoring is stronger, $\gamma\geq 0.5$, the boundary region is much smaller than subsystem size $\ell\geq 16$ that we investigate, so $P(\Delta S_{qj})$ do not depend on the subsystem size. For the projective measurement protocol, we also obtain similar results (not shown).
@@ IV.2 Spatial dependence of $\Delta S_{qj}$
An explicit calculation of the distribution function separating sites close to the boundary from those in the bulk, see Fig. 7, confirms the qualitative picture above. For weak monitoring, the distribution function for the boundary point is a Gaussian with a small peak at zero. As more sites around the boundary are added, the distribution develops a narrow exponential tail for $\Delta S_{qj}\gtrsim 0$. We note that rare instances of $\Delta S_{qj}\gtrsim 0$ occur due to the fact that the measurement not only disentangles the target site but also rearranges the remaining particles across the rest of the L1L-1 sites, effectively enhancing entanglement in unmeasured regions. This counterintuitive effect highlights the non-locality inherent in quantum measurement dynamics. For bulk sites, the distribution is asymmetric with exponential tails characterized by different decay rates. For strong monitoring, the distribution is effectively a delta function at zero for bulk points (note the scale of the axis), while for boundary points, though there is also a peak at zero, we observe an asymmetric distribution with a broad support that covers all possible values of $\Delta S_{qj}$.
@@
Figure 7: The distribution function of $\Delta S_{qj}$ in Eq. (9) for specific sites. The upper panel: the monitoring rates is $\gamma=0.1$ and the system size is $L=1024$. The bottom panel: the monitoring rates is $\gamma=3.0$ and the system size is $L=512$. Here sites 1, 2 and 3 stand for sites around the boundary of the subsystem. Site 4 stands for the site at the center of the subsystem. Specifically, sites 1 are sites $[r=1,r=L/2,r=L/2+1,r=L]$, sites 2 are [r=1+5,r=L/25,r=L/2+6,r=L5][r=1+5,r=L/2-5,r=L/2+6,r=L-5], sites 3 are [r=1+15,r=L/215,r=L/2+16,r=L15][r=1+15,r=L/2-15,r=L/2+16,r=L-15] and sites 4 $[r=L/4,r=3L/4]$. Note that for Fig. 7, the scale of $\Delta S_{qj}$ is of the order $10^{-10}$ and the maximum is $10^{10}$ so it is effectively a Dirac delta function.
Figure 8: The density-map of $\Delta S_{qj}$ in Eq. (9) as a function of the occupation number of the selected site that are going to be measured. The color scale indicates the normalized density of measurement events. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 8 to 8. The system size for the upper panel is $L=1024$, and the lower panel is $L=512$. The red dashed line is Eq.(10), signals the maximum reduction of the entanglement entropy after a single quantum-jump measurement.
IV.3 Effects of the occupation number $n_{i}$ on $\Delta S_{qj}$
@@
In the strong monitoring limit, even though $n_{i}$ can take any value within its domain of definition $[0,1]$, the change in EE has a maximum value of $\ln{2}\approx 0.69$, reflecting the maximal entanglement contribution from a single particle. As mentioned earlier, this maximum value can only occur at the subsystem boundary sites because, for the rest of sites, measurements induce little change in $\Delta S_{qj}$. Therefore, a measurement would only affect particles around the boundary. In this strong coupling limit, it is possible to compute analytically the maximum decrease of the entanglement entropy after each quantum jump measurement. For that purpose, we assume that after each measurement, only the measured particle is affected, while the rest of the system remains unchanged. Therefore, the measurement only affects the entropy associated with site $i$, with no rearrangement effects that could increase the total EE. We have to distinguish two cases, if the occupation number of the site $i$ before the measurement is $n_{i}>0.5$, the maximum bipartite entanglement entropy is dominated by the entanglement between this site $i$ and the other half subsystem, which is given by $S=-n_{i}\ln{n_{i}}-(1-n_{i})\ln{(1-n_{i})}$. If $n_{i}<0.5$, the maximum entanglement is no longer dominated by the entanglement between the site $i$ and the other half system because the particle wave function can be extended over several sites across the subsystem boundary, namely, effectively the particle is fully entangled only in the subsystem boundary, leading to a maximum single-particle entanglement entropy $S=\ln{2}$. After measurement, the entanglement entropy of this particle drops to 0. Therefore, the maximum drop in entanglement entropy induced by a single quantum-jump measurement is then given by
$\Delta S_{qj}=\begin{cases}-\ln{2}&n_{i}<0.5\\ n_{i}\ln{n_{i}}+(1-n_{i})\ln{(1-n_{i})}&n_{i}\geq 0.5\end{cases}$ (10)
This assumption shows excellent agreement with our numerical results, as illustrated in Fig. 8. In contrast, the maximum increase of entanglement entropy after a measurement, which also shows a sharp bound, remains less understood, because it requires a coordinated rearrangement of many fermions near the subsystem boundary so it cannot be expressed solely in terms of $n_{i}$. A detailed characterization of this upper bound is beyond the scope of this paper.
IV.4 The recovery of entanglement due to non-Hermitian evolution
@@
where $\tau$ is the time between quantum jumps and $\Delta S_{nH}=S_{A}(t+\tau)-S_{A}(t)$ computed from Eq. (7). We use $\delta$ in $\delta S_{nH}$ to distinguish the time dimension from $\Delta$ in $\Delta S_{nH}$, consistent with $\delta S_{qsd}$. Therefore, $\delta S_{nH}$ captures the system intrinsic capacity to regenerate entanglement between measurements. Since we are interested in the late time dynamics of entanglement for which the average EE has reached the saturation value and therefore does not have a net growth in time, we expect [68] that, on average, the generation of entanglement due to the non-Hermitian evolution will compensate its destruction due to measurements $\overline{\delta S_{nH}}\approx\overline{\Delta S_{qj}}/\bar{\tau}$, where $\overline{\penalty 10000\ \bullet\penalty 10000\ }$ stands for the average over quantum trajectories. Results depicted in the left plot of Fig. 9 confirm that this is the case for all considered monitoring strengths. We stress that this cancellation is a property of the average only. We shall see that the distribution of $\delta S_{nH}$ and other properties are qualitatively different from those due to quantum jumps $\Delta S_{qj}$.
@@
Figure 9: The average change in EE after a measurement as a function of the monitoring rate $\gamma$ using the 9 QJ protocol (left) and the 9 PM protocol (right). This is compared with the average speed of entanglement gain between measurements in the QJ protocol ($\delta S_{nH}$) (left) and the PM protocol ($\delta S_{un}$) (right). As is expected, the creation of entanglement is compensated by the loss due to the measurement. The system size is $L=512$.
@@
Figure 10: The probability distribution of the change of EE $\delta S_{nH}$ in Eq. (11), computed using Eq. (7), between measurements due to the non-hermitian evolution in the QJ protocol for different system sizes and times longer than the saturation time. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 10 to 10. The monitoring rate is the same as in Fig. 4.
IV.5 Probability distribution of $\delta S_{nH}$
@@
Figure 11: The entanglement change $\delta S_{nH}$ in Eq. (11) due to the non-Hermitian evolution in the QJ protocol after measuring site $r$. The color scale indicates the normalized density of measurement events. The left plot is for the weak monitoring rate $\gamma=0.1$ and system size $L=1024$, while the right plot is for the strong monitoring rate $\gamma=3$. The system size is $L=512$. Different from the effect of quantum jumps $\Delta S_{qj}$, see Fig. 5, $\delta S_{nH}$ due to the non-Hermitian evolution does not show any dependence on the measuring site.
Figure 12: The density-map of $\delta S_{nH}$ in Eq. (11) for the QJ protocol as a function of the occupation number before the measurements. The color scale indicates the normalized density of measurement events. The monitoring rate are $\gamma=0.1$ (left), and $\gamma=3.0$ (right).
IV.6 Site dependence of $\delta S_{nH}$
@@ IV.7 The scaling behavior of the mutual information $I(d_{r})$
Figure 13: The mutual information $I(d_{r})$ in Eq. (12) between the subsystem AA and a site $M_{r}$ in the other subsystem as a function of the distance $d_{r}$ for the QJ protocol. The pink dashed line is a power-law fitting and the red dashed line is an exponential fitting. The system size is $L=512$.
@@ V The projective measurement protocol
Finally, we study the PM protocol corresponding to the standard picture in quantum mechanics where, after a measurement, the wavefunction collapses to one of the eigenstates of the Hermitian operator representing the observable being measured with a probability given by the Born’s rule. Details of the protocol are given in Appendix A.3. This protocol have similarities with the quantum jump one but with three differences. First, in the projective measurement protocol, at each time step, the probability of measuring each site is the same while in the quantum jump protocol it depends on the result of the previous measurement. This is a consequence of the fact that only the quantum jump is a continuous measurement protocol. In the projective measurement protocol, the evolution between measurements is unitary. Second, the outcome of the quantum jump protocol is only $n_{i}=1$, while the outcome of the projective measurement is either $n_{i}=0$ or $n_{i}=1$. As a consequence, in the strong monitoring limit, we expect quantum Zeno effect around these two values. Third, the waiting time $\tau$ between measurements can be chosen freely in the projective measurement protocol while in the quantum jump protocol is dictated by the non-Hermitian decay. Despite these differences, we show next that the results is very similar to that of the quantum jump protocol.
Figure 14: The probability distribution $P(\Delta S_{pm})$ of the entanglement entropy change $\Delta S_{pm}$ in Eq. (13), computed from Eq. (7) after saturation of the EE, due to a projective measurement of the occupation number (PM protocol). For convenience, the distribution $P(\Delta S_{pm})$ is multiplied by the system size $L$. The monitoring rates are $\gamma=0.1,0.5,1.5$ and $3.0$ from 14 to 14.
Figure 15: Density map of $\Delta S_{pm}$ in Eq. (13) with the same parameters as in Fig. 8 (QJ protocol) but for the PM protocol. The color scale indicates the normalized density of measurement events. Results are similar for both protocols after taking into account that the potential outcomes in the PM protocol are $n_{i}=0,1$ while in the QJ protocol is only $n_{i}=1$.
As in the previous cases, we also define the change in EE
@@
Figure 16: The density-map of $\delta S_{un}$ in Eq. (13) for the PM protocol as a function of the occupation number right before the measurement. The color scale indicates the normalized density of measurement events. The monitoring rates are $\gamma=0.1$ (left), and $\gamma=3.0$ (right).
@@ VI Discussion and Conclusions
We have studied the distribution function of the change in EE after a measurement of the occupation number in a chain of free fermions using the QSD, QJ and PM protocols. The distribution for the QSD protocol is Gaussian and size independent in the weak monitoring limit. As the monitoring strength increases, the distribution is still symmetric but deviations from Gaussianity are clearly observed which are well characterized by a distribution which is Gaussian around zero but with exponential tails. In the strong monitoring limit, the tail are still exponential but the Gaussian around zero turns into a strong peak suggesting a singularity at zero which is a signature of the Zeno effect.
QJ and PM protocols lead to asymmetric non-Gaussian distributions with a peak at zero that becomes stronger as the monitoring strength or the system size increases. More interestingly, for these two protocols, the distribution function is spatially inhomogeneous, namely, it is not the same for all sites. For boundary sites, separating the two sub-systems in the definition of the EE, the distribution in the weak monitoring limit is close to a Gaussian with a small peak at zero and a broad support. However, for the rest of sites, the support is one order of magnitude smaller and the distribution has asymmetric exponential tails. On average, the change in EE is negative, however, we have found that the distribution has an exponential tail for positive values so there is a finite probability that the EE increases as a result of a single measurement. It would be interesting a detailed characterization of the conditions for the existence of such measurements resulting in a positive change of the EE.